# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ A.5 Uniform (All-$\mathbf{x}$) Bound on the Parallel Term
Bounding the Approximation Error: We seek to understand the gap between $\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ and $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ For any $\mathbf{x}\in\mathcal{K}$ we let $\mathbf{x}_{r}=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$. Here, $\mathbf{x}_{r}$ depends on $\mathbf{x}$, but we drop such dependence to avoid cumbersome notation. Note that $\|\mathbf{x}-\mathbf{x}_{r}\|_{2}\leq r$, and indeed we have $\mathbf{x}-\mathbf{x}_{r}\in\mathcal{K}_{(r)}=(\mathcal{K}-\mathcal{K})\cap\mathbb{B}_{2}^{n}(r).$ For clarity we consider a given $\mathbf{x}\in\mathcal{K}$, while we note beforehand that the final bounds in (A.30) and (A.33) hold uniformly for all $\mathbf{x}\in\mathcal{K}$. Now by $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ defined in (A.15), we calculate that
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}-\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big{|}\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})-f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big{|}$ (A.27)
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}\langle\mathbf{x}_{0}-\mathbf{x}_{r},\mathbf{u}\rangle\big{|}$
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}+r.$ (A.28)
@@
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}$
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}$
$+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{|}:=I_{1}+I_{2}.$ (A.29)
@@
$I_{1}$ $\leq\frac{1}{\kappa m}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\big{|}\bm{\Phi}_{i}^{*}\mathbf{u}\big{|}^{2}\right)^{1/2}$
$\leq\frac{2\eta}{\kappa}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq\eta m\end{subarray}}\left(\frac{1}{\eta m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}\leq C_{4}\eta\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{\eta m}}+\sqrt{\log(\eta^{-1})}\right),$ (A.30)
@@
$I_{2}\leq\frac{2}{\kappa\eta}\frac{1}{m}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\big{|}\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})\big{|}\big{|}\bm{\Phi}_{i}^{*}\mathbf{u}\big{|}$ (A.31)
$\leq\frac{2}{\kappa\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$ (A.32)
@@
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\right|\leq C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.34)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}$ $\leq\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}+r$
$+C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right),$
and then we further apply (A.25), as well as $\eta<1$, to arrive at
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{|}f_{1}^{\|}(\mathbf{x},\mathbf{u})\big{|}\leq C_{8}\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{\frac{\mathscr{H}(\mathcal{K},r)}{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.35)
In summary, this bound holds with probability at least $1-C_{9}\exp(-c_{10}\eta m)-C_{11}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathcal{H}(\mathcal{K},r))$ under the sample complexity $m=\Omega(\frac{\mathscr{H}(\mathcal{K},r)}{\eta}+\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see such sample complexity and probability term from (A.26), (A.25), (A.30) and (A.29). We now set $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ as in our statement, and the bound in (A.35) reads as
$O\left(\frac{\omega(\mathcal{U}_{1}^{\rm(N)})}{\sqrt{m}}+\sqrt{\frac{\mathscr{H}(\mathcal{K},r)}{m}}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}\right).$ (A.36)
Under the assumed sample complexity (A.24), the bound (A.36) scales as $O(\eta\sqrt{\log(\eta^{-1})})$, and we can promise a probability at least $1-C_{12}\exp(-\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$ for the claim to hold. This completes the proof. ∎
@@ A.6 Uniform (All-$\mathbf{x}$) Bound on the Orthogonal Term
In the setting of Lemma A.5, there exist some absolute constants $c_{1},C_{2},C_{3},c_{4},C_{5},c_{6},C_{7}$ such that given any $\eta\in(0,c_{1})$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if
$m\geq C_{2}\left(\frac{\omega^{2}(\mathcal{U}_{c}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}\right),$ (A.37)
then with probability at least $1-C_{3}\exp(-c_{4}[\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},\eta^{3})])-C_{5}\exp(-c_{6}\eta^{2}m)$, we have
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq C_{7}\eta\sqrt{\log(\eta^{-1})}.$
@@
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{1}(\omega(\mathcal{U}_{c}^{\rm(N)})+t)}{\sqrt{m}}$
holds with probability at least $1-2\exp(\mathscr{H}(\mathcal{K},r)-t^{2})$. Therefore, setting $t^{2}=\Theta(\omega^{2}(\mathcal{U}_{c}^{\rm(N)})+\mathscr{H}(\mathcal{K},r))$ yields that the event
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}\leq\frac{C_{2}(\omega(\mathcal{U}_{c}^{\rm(N)})+\sqrt{\mathscr{H}(\mathcal{K},r)})}{\sqrt{m}}$ (A.38)
holds with probability at least $1-\exp(-\omega^{2}(\mathcal{U}_{c}^{\rm(N)})-\mathscr{H}(\mathcal{K},r))$.
@@
$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big{|}-\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$ (A.39)
$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\big{|}f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})-f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big{|}$
$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}-\Big{\|}\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}\Big{|}$ (A.40)
$\quad\quad\quad\quad+\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\Big{\|}\begin{bmatrix}\mathbf{u}_{\mathbf{x}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big{\|}_{2}-\Big{\|}\begin{bmatrix}\mathbf{u}_{\mathbf{x}_{r}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big{\|}_{2}\Big{|}.$ (A.41)
@@
By the triangle inequality and the observation that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ gives $\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}$, the term in (A.40) is bounded by
@@
$=\frac{1}{\sqrt{m}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{\|}\Im\big{[}\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big{]}\big{\|}_{2}.$ (A.42)

### v3

@@ A.5 Uniform (All-$\mathbf{x}$) Bound on the Parallel Term
Bounding the Approximation Error: We seek to understand the gap between $\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ and $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ For any $\mathbf{x}\in\mathcal{K}$ we let $\mathbf{x}_{r}=\mathrm{arg}\min_{\mathbf{u}\in\mathcal{N}_{r}}\|\mathbf{u}-\mathbf{x}\|_{2}$. Here, $\mathbf{x}_{r}$ depends on $\mathbf{x}$, but we drop such dependence to avoid cumbersome notation. Note that $\|\mathbf{x}-\mathbf{x}_{r}\|_{2}\leq r$, and indeed we have $\mathbf{x}-\mathbf{x}_{r}\in\mathcal{K}_{(r)}=(\mathcal{K}-\mathcal{K})\cap\mathbb{B}_{2}^{n}(r).$ For clarity we consider a given $\mathbf{x}\in\mathcal{K}$, while we note beforehand that the final bounds in (A.30) and (A.33) hold uniformly for all $\mathbf{x}\in\mathcal{K}$. Now by $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ defined in (A.15), we calculate that
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|-\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big|\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})-f_{1}^{\|}(\mathbf{x}_{r},\mathbf{u})\big|$ (A.27)
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|\langle\mathbf{x}_{0}-\mathbf{x}_{r},\mathbf{u}\rangle\big|$
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|+r.$ (A.28)
@@
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|$
$\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|$
$+\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big|:=I_{1}+I_{2}.$ (A.29)
@@
$I_{1}$ $\leq\frac{1}{\kappa m}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}|\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\sum_{i\in\mathcal{J}_{\mathbf{x}_{r},\eta}}\big|\bm{\Phi}_{i}^{*}\mathbf{u}\big|^{2}\right)^{1/2}$
$\leq\frac{2\eta}{\kappa}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|\leq\eta m\end{subarray}}\left(\frac{1}{\eta m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}\leq C_{4}\eta\left(\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{\eta m}}+\sqrt{\log(\eta^{-1})}\right),$ (A.30)
@@
$I_{2}\leq\frac{2}{\kappa\eta}\frac{1}{m}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\sum_{i\notin\mathcal{J}_{\mathbf{x}_{r},\eta}}\big|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})\big|\big|\bm{\Phi}_{i}^{*}\mathbf{u}\big|$ (A.31)
$\leq\frac{2}{\kappa\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}\right)^{1/2}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$ (A.32)
@@
$\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\bm{\Phi}_{i}^{*}\mathbf{u}\big)\right|$
$\leq C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.34)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|$ $\leq\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|+r+C_{7}\left(\frac{\sqrt{\eta}\cdot\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}+\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right),$
and then we further apply (A.25), as well as $\eta<1$ and $\frac{\omega(\mathcal{U}_{1}^{\mathbb{S}})}{\sqrt{m}}\leq\eta\sqrt{\log(\eta^{-1})}$ that holds under $m=\Omega(\frac{\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})})$, to arrive at
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big|f_{1}^{\|}(\mathbf{x},\mathbf{u})\big|\leq C_{8}\left(\eta\sqrt{\log(\eta^{-1})}+\frac{\omega(\mathcal{K}_{(r)})}{\eta\sqrt{m}}+\frac{r}{\eta}\right).$ (A.35)
In summary, this bound holds with probability at least $1-C_{9}\exp(-c_{10}\eta^{2}\log(\eta^{-1})m)$ under the sample complexity $m=\Omega(\frac{1}{\eta^{2}\log(\eta^{-1})}[\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}})+\mathscr{H}(\mathcal{K},r)])$; see such sample complexity and probability term from (A.25), (A.26) and (A.30). We now set $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$, and using the assumed sample complexity (A.24), then the bound in (A.35) reads as $O(\eta\sqrt{\log(\eta^{-1})})$. This completes the proof. ∎
@@ A.6 Uniform (All-$\mathbf{x}$) Bound on the Orthogonal Term
In the setting of Lemma A.5, there exist some absolute constants $c_{1},C_{2},C_{3},c_{4},C_{5}$ such that given any $\eta\in(0,c_{1})$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if
$m\geq C_{2}\left(\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}\right),$ (A.36)
then with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}m)$, we have
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|{f}_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|\leq C_{5}\eta\sqrt{\log(\eta^{-1})}.$
@@
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\frac{C_{1}(\omega(\mathcal{U}_{c}^{\mathbb{S}})+t)}{\sqrt{m}}$
holds with probability at least $1-2\exp(\mathscr{H}(\mathcal{K},r)-t^{2})$. Therefore, under the sample complexity $m=\Omega(\frac{1}{\eta\log(\eta^{-1})}[\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})+\mathscr{H}(\mathcal{K},r)])$, setting $t=\sqrt{\eta\log(\eta^{-1})m}$ yields that the event
$\sup_{\mathbf{x}\in\mathcal{N}_{r}}\sup_{\mathbf{u}\in\mathcal{U}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|\leq C_{2}\sqrt{\eta\log(\eta^{-1})}$ (A.37)
holds with probability at least $1-2\exp(-\frac{1}{2}\eta\log(\eta^{-1})m)$.
@@
$\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})\big|-\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big|$ (A.38)
$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\big|f_{1}^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})-f_{1}^{\bot}(\mathbf{x}_{r},\mathbf{u},\mathbf{w})\big|$
$\leq\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}-\Big\|\frac{\Im(\operatorname{diag}(\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r}))^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}\Big|$ (A.39)
$\quad\quad\quad\quad+\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\Big\|\begin{bmatrix}\mathbf{u}_{\mathbf{x}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big\|_{2}-\Big\|\begin{bmatrix}\mathbf{u}_{\mathbf{x}_{r}}^{\bot}\\ \mathbf{w}\end{bmatrix}\Big\|_{2}\Big|.$ (A.40)
@@
By the triangle inequality and the observation that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$ gives $\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}$, the term in (A.39) is bounded by
@@
$=\frac{1}{\sqrt{m}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big\|\Im\big[\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big]\big\|_{2}.$ (A.41)